OpenStudy (yanasidlinskiy):

Is this right? The velocity of sound in air is given by the equation =20√(273+t) , where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity is 329 meters per second by graphing the equation. Round the answer to the nearest degree.

OpenStudy (astrophysics):

\[329 = 20\sqrt{273+t}\] Solve for t.

ganeshie8 (ganeshie8):

congrats on reaching \(\large \color{gold}{75!} \)

OpenStudy (yanasidlinskiy):

\[\frac{ 329 }{ z } = \frac{ 10\sqrt{273+t} }{20 }\] \[(6.45)^2 = (\sqrt{273 + t} )^2\] \[279.6 = 273 + t\] -273 -273 t = -2

OpenStudy (astrophysics):

Yes, actually that works to.

OpenStudy (yanasidlinskiy):

Hahahha, Lol. Thanx @ganeshie8 I was sooo surprised when I've reached it:D!! Check my answer:D What did I do wrong?

OpenStudy (astrophysics):

Good job @YanaSidlinskiy and congrats on 75 xD

OpenStudy (yanasidlinskiy):

Thanx @Astrophysics yall are the best!!:D You too ganeshie:D!!! Great on that LaTeX--> 75:D

OpenStudy (yanasidlinskiy):

I have another one. I need serious help though:D

OpenStudy (neer2890):

but i think you asked for graphing the equation.???

OpenStudy (astrophysics):

That should be cake with t :)

OpenStudy (yanasidlinskiy):

On my homework, it doesn't have a graph anyways. The main thing I want for this is ya'll to check if I have it set up and solved correctly!

OpenStudy (yanasidlinskiy):

Lol, Astrophics:D